# Ring Deconvolution Microscopy: An Exact Solution for Spatially-Varying Aberration Correction

### v3

@@ Abstract
The most ubiquitous form of computational aberration correction for microscopy is deconvolution. However, deconvolution relies on the assumption that the point spread function is the same across the entire field-of-view. It is well recognized that this assumption is often inadequate, but space-variant deblurring techniques generally require impractical amounts of calibration and computation. We present a new imaging pipeline, ring deconvolution microscopy (RDM), that leverages the rotational symmetry of most optical systems to provide simple and fast spatially-varying aberration correction. We formally derive theory and algorithms for exact image recovery and additionally propose a neural network based on Seidel coefficients as a fast alternative. We showcase significant enhancements both visually and quantitatively compared to standard deconvolution and existing spatially-varying deconvolution across a diverse range of microscope modalities, including miniature microscopy, multicolor fluorescence microscopy, and point-scanning multimode fiber micro-endoscopy. Our approach enables near-isotropic, subcellular resolution in each of these applications.
@@
Figure 1: Ring Deconvolution Microscopy (RDM). a) Point sources at the sample plane (left) are imaged (right) to point spread functions (PSFs) with a rotationally symmetric imaging system. The PSFs are linear revolution-invariant (LRI)—they vary with distance from the center of the field-of-view (FoV) (top row), but maintain the same shape at a fixed radius $r$, just revolved around the center (bottom row). b) The RDM pipeline. A one-time calibration procedure (top) captures a single image of randomly-placed point sources (e.g., fluorescent beads), then fits the primary Seidel coefficients (see Sec. 4.2). Next, we either use the Seidel coefficients to generate a radial line of synthetic PSFs, if using ring deconvolution, or we feed the coefficients directly into Deep Ring Deconvolution (DeepRD) . After calibration, we can deblur images (bottom) using either ring deconvolution or DeepRD. c) Experimental deblurring of live tardigrade samples imaged with the UCLA Miniscope [1]. From left to right: measurement, standard deconvolution, ring deconvolution, and DeepRD. Ring deconvolution and DeepRD consistently outperform deconvolution.
@@ 1 Introduction
Much of optical engineering is focused on reducing aberrations by adding additional corrective optical elements to an imaging system—consider a microscope objective, comprised of numerous lenses stacked in a housing. While such designs allow for high-performance imaging, they also incur added cost, weight, and complication. Even with very large and expensive lens stacks, it is difficult and, in some cases, impossible to correct all aberrations across a large area, and so aberrations are often what limit the usable field-of-view (FoV) of a system. Furthermore, some systems cannot accommodate any aberration-correction optics; for example, additional elements may not fit in miniaturized microscopes (e.g. Miniscope [1]) and are prohibitively expensive for large-aperture telescopes [2].
@@
Faced with a poorly-corrected imaging system, the modern microscopist instead turns to computational aberration correction, where the burden is shifted onto computer algorithms applied after the image is acquired. The most commonly used correction technique, image deconvolution, captures a calibration image of a small point-like source, known as the point spread function (PSF), in order to characterize the aberrations. The PSF can then be used to computationally deconvolve any image taken with the system via simple and fast algorithms, to yield a deblurred result. A main limitation of this approach is that it assumes that the system’s PSF does not vary spatially (i.e. the system is linear space-invariant, LSI). This assumption is usually only true near the center of the FoV, and optical designers often artificially sacrifice part of the system’s FoV in order to maintain LSI.
@@
To bypass the space-invariant limitations, a large community effort has gone towards heuristic forms of spatially-varying ’deconvolution’, wherein one measures PSFs at multiple points in the FoV and uses them to correct the image. Such heuristics include: assuming each region of an image is locally LSI [3], adaptively splitting the FoV by first quantifying the degree of space-variance [4, 5], interpolating PSFs [6, 7], decomposing the PSF into space-invariant orthogonal modes [8, 9, 10, 11, 12, 13, 14, 15], and doing the same in Fourier space [16]. As the number of PSFs collected grows—into the hundreds of thousands—these heuristics approach exact recovery. However, the tradeoff in terms of the complexity of calibration and computation quickly becomes intractable. For example, in patch-wise deconvolution, the FoV is divided into patches, each of which gets deconvolved by a PSF measured at its center. The maximum accuracy is achieved when the patch size is reduced down to a single pixel, but even for a single 1 megapixel image, deblurring would require a million PSF measurements and a computation time of hundreds of hours (see Fig. 5c).
@@
Another emerging modality is deep deblurring [17, 18, 19, 20, 21, 22, 23] , in which varying amounts of system information are incorporated into a deep neural network. On one end of the spectrum, networks that are primarily data-driven struggle with extrapolation beyond the training data, and tend to reproduce whatever biases existed therein, a particularly relevant point since many of them are trained on simulated data. On the other end, networks that incorporate physical information, such as PSFs, may have better generalization properties but suffer the same accuracy/efficiency tradeoff as patch-based methods. For these reasons, spatially-varying deblurring has not become commonplace among practitioners. Thus, there is a need for a new method of calibrating and deblurring spatially-varying imaging systems which is effective, efficient, and robust.
Here, we propose a spatially-varying method that requires only a single calibration image and has reasonable compute time, while offering theoretically exact deblurring for imaging systems that are rotationally symmetric—i.e., symmetric about their optical axis. This symmetry occurs in many imaging systems by design, and a significant portion of optical theory is developed under this assumption. While some existing deblurring techniques have leveraged this symmetry for modeling motion blur [24, 25], for suppressing radial variance [26], and for deblurring sections of the FoV in space-invariant radial segments [27, 28, 29, 30, 31].
@@
Our technique, Ring Deconvolution Microscopy (RDM), is a new imaging pipeline that achieves both accuracy and efficiency. We rigorously prove that RDM enables exact recovery for a rotationally-symmetric imaging system, but it is also practical, both computationally and in terms of calibration. The first step in RDM is a simple, single-shot calibration scheme, wherein the system’s primary Seidel aberration coefficients are estimated from a single image of randomly-distributed point sources. These coefficients quantify the severity of spatial variance and provide the necessary system information for the second step, deblurring. Within this second step there are two proposed image deblurring algorithms. Our first and main algorithm, ring deconvolution, uses a new and exact theory for rotationally symmetric imaging to deblur the image at all points in the FoV, with only order $N^{3}\log(N)$ ($N$ is the image side length) compute time, as compared to $N^{4}$ for full spatially-varying deblurring. For even faster computation, but without guarantees on out-of-distribution accuracy, we further propose an alternate neural network-based algorithm called Deep Ring Deconvolution (DeepRD), which constrains learning with physical knowledge provided by the system’s Seidel aberration coefficients.
@@
These algorithms outperform existing methods, approaching subcellular, isotropic resolution across the FoV. We demonstrate this on three microscope modalities: miniature microscopy, multicolor fluorescence microscopy, and multimode fiber micro-endoscopy. See Fig. 1 for a summary of the pipeline along with an example result on images of live tardigrades. An open source implementation of RDM can be found here.
@@ 2 Results
2.1 Single-shot calibration
The system is calibrated by measuring its response to a point source, i.e., its PSF. The PSFs of space-varying systems will be different at different parts of the FoV, and many PSF measurements may be required to fully characterize the system. RDM leverages the fact that rotationally symmetric imaging systems require fewer measurements for exact system characterization.
The reduction in complexity occurs because PSFs that are the same distance from the center of the FoV will all be the same shape, but rotated at different angles (see Figure 1a). We call this property of the PSFs linear revolution-invariance (LRI), and denote it mathematically as
@@ 2.1 Single-shot calibration
where $\tilde{h}(\rho,\phi;r,\theta)$ is the (spatially-varying) PSF in polar coordinates from a point source at location $(r,\theta)$. Note that the shape of the PSF itself is not necessarily rotationally symmetric. Intuitively, each point along a concentric ring in the FoV will have the same PSF shape, and the system can be exactly characterized by measuring the PSF only once for each radius $r$ from the optical center.
In practice, directly measuring the PSF at every radius $r$ is impractical. Instead, under LRI, there is a simple and effective method for simultaneously estimating these PSFs from a single image. The method is described in Section 4.2 and is the first step of the RDM pipeline. It works by estimating the Seidel aberration coefficients from an image of point sources randomly scattered across the FoV (see Fig. 1b). No motion stage is needed. The Seidel aberrations are similar to the Zernike modes [32, 33] but much simpler to estimate—Seidel polynomials are explicitly functions of the field position and thus use the same coefficients regardless of the position in the FoV [34]. Meanwhile, Zernike modes are a general-purpose spherical basis set, which is not a function of field position and whose coefficients are therefore different at each field position. Our experiments demonstrate that the 5 primary (3rd order) coefficients (commonly called sphere, coma, astigmatism, field curvature, and distortion) suffice to characterize spatially-varying LRI systems.
2.2 Reconstruction algorithm
After calibration, the next step is to use the estimated Seidel coefficients to algorithmically deblur the measured image. LRI systems, much like LSI systems, allow for computationally efficient and exact forward and inverse models. Section 4.1 provides a description of these algorithms, which we call ring convolution and ring deconvolution, respectively. Therein is also included a theoretical proof of their exactness. We briefly describe ring deconvolution here as our method for aberration correction, along with DeepRD, its extension to deep learning. These methods form the second step in the RDM pipeline (Fig. 1b).
@@ 2.2 Reconstruction algorithm
    Ring deconvolution. We derive a provably exact algorithm for reconstructing the underlying sample from a blurry image given the PSF at each radius of the FoV of a rotationally symmetric imaging system. These PSFs are efficiently obtained via the single-shot calibration procedure described above. This is our main algorithm.
@@
    Deep Ring Deconvolution (DeepRD). Although ring deconvolution is significantly faster to compute than a full patch-based spatially-varying deblur technique, it may still be relatively slow (on the order of a few minutes) for very large image sizes (e.g., beyond 1024×1024)1024$\times$ 1024)1024 × 1024 ) or video data. We used deep learning to perform a fast (but approximate) version of ring deconvolution called DeepRD. As input, it takes a blurry image and a list of the 5 primary Seidel coefficients. DeepRD is trained on a dataset of natural images synthetically blurred using ring convolution. The experimental images are completely out of the training domain, yet the algorithm generalizes effectively.
Even when the system is LSI or close to LSI, RDM still offers an improvement over standard deconvolution by instead deconvolving with a synthetic PSF generated by the Seidel coefficients (see Sec. 2.7).
@@
Figure 2: RDM on miniature microscopy. After calibrating the Miniscope with a single image of fluorescent beads (see Fig. 1b), we estimate the Seidel coefficients. We then show the measurement and results from several deblurring methods for comparison: standard deconvolution (using a single measured PSF), a U-Net trained on our spatially-varying blur dataset, and our methods: ring deconvolution and DeepRD. Deconvolution assumes space-invariance while the remaining methods are designed to handle spatially-varying aberrations. In the first row, ring deconvolution and DeepRD can resolve resolution target elements near the edges of the FoV, which are not resolved by the other methods. Zoom-ins show RDM resolves up to element 6 of group 9 (blue inset) and element 5 of group 8 (green inset). In the middle row, both RDM methods more sharply resolve membranes in the rabbit liver tissue (blue inset) and a reveal a hook-shaped feature not seen using the other methods (green inset). In the bottom row, our methods are also able to resolve features on the tardigrade not well resolved by the other methods (yellow inset) and improve the contrast compared to deconvolution throughout the image (green and blue insets).
2.3 Miniature microscopy
@@
Our first experimental system is the UCLA Miniscope [1], a miniature microscope used largely for neuro-imaging in freely-behaving animals. The first step of our experimental pipeline is to calibrate the imaging system by capturing a single image of randomly-placed fluorescent beads (obtained by smearing bead solution on a slide). We use it to fit Seidel coefficients, obtaining the coefficients 0.85, 0.56, 0.25, 0.29, and 0 waves of spherical aberration, coma, astigmatism, field curvature and distortion, respectively. These numbers, while specific to our particular assembly of the Miniscope, are consistent with the aberration profile of a radial GRIN lens [35], which is the objective lens used by the Miniscope. Radial GRIN lenses are dominated mostly by spherical aberration followed by coma, astigmatism, and field curvature. The fact that the off-axis coefficients (i.e. all of the primary coefficients except for spherical) are nonzero confirms that the system is indeed spatially-varying. For comparison, we also run the standard deconvolution calibration, in which a single fluorescent bead is isolated and centered in the FoV, thus allowing us to measure only the center PSF. The PSF is then denoised before its use in deconvolution (see 4.5).
@@
Having completed the one-time calibration for the Miniscope, we image a variety of different samples: a USAF resolution target, rabbit liver tissue, and live fluorescence-stained tardigrades. For each sample, we run ring deconvolution and DeepRD, and show a comparison with standard deconvolution (using a single PSF measured at the center of the FoV) and a baseline U-Net (see Fig. 2).
@@
The USAF resolution target was placed at 9 separate locations in the FoV, in order to observe the smallest features at each location. We stitch these 9 images together, using only the region of each constituent image that contains the high resolution group, to form the measurement image shown in Fig. 2 (first row). Ring deconvolution and DeepRD—having knowledge of the field-varying aberrations via the Seidel coefficients—give the most improvement near the edges and corners of the image. The standard deconvolution produces a noisy, low contrast result in those regions due to the mismatch of the center PSF with the edge PSFs. We also note that both of the learning methods (U-Net and DeepRD) offer the best denoising performance, a well-known property of neural networks [36]; however, this comes at the cost of inconsistent performance—both models perform worse on the resolution target than on the other samples.
@@
For our second sample, rabbit liver tissue, the RDM methods perform well, revealing features in the corners of the image including the outlines of membranes, which are not clear in the raw measurement or after the other methods. Such clarity is critical for downstream tasks like image segmentation.
@@
Finally, we fluorescently stain live, semi-starved tardigrades with a DNA gel stain, and capture a series of videos. We apply deblurring to each frame and display one such frame in the bottom row of Fig. 2; the full videos can be found here. As with the previous experiments, ring deconvolution and DeepRD provide increased image contrast and detail in the corners of the frames as compared to the other methods. In particular, the small, dot-like features within the tardigrade are better resolved.
@@
Further experimental details can be found in Sec. 4.5. It is worth noting that, though the baseline U-Net we use is inspired by the popular Content-Aware Image Restoration (CARE) model [21], we train both DeepRD and the baseline U-Net using a novel synthetic training strategy: applying ring convolutions using PSFs generated from random Seidel coefficients to natural images from Div2k [37](see Sec. 4.3). We will release this synthetic dataset upon publication. Despite being trained on simulated data, both the U-Net and DeepRD generalize surprisingly well to the images taken with the Miniscope. Further, in Appendix A, we provide a demonstration that DeepRD is interpretable: it understands the individual effects of each aberration coefficient.
In summary, We find that ring deconvolution is consistently the best method, a result that reflects the fact that it is theoretically exact for rotationally symmetric systems and is not biased toward any particular distribution of images. DeepRD performs similarly to ring deconvolution, but it is faster and less consistent. Both perform better than the U-Net and standard deconvolution.
Figure 3: RDM on multicolor fluorescence imaging. a) Fluorescent beads imaged with a 100x 1.4 NA objective (top) and b) corresponding Seidel-fitted PSFs (bottom). The system is spatially-varying; PSFs in the corner are larger and elongated (blue inset) compared to the center PSF (yellow inset). Seidel-fitting faithfully capture this. c) Two examples of BPAE cells (left), processed by standard deconvolution (middle), and RDM (right). Top row: Deconvolution and RDM perform similarly in the center (yellow inset) but RDM is better in the corner (blue inset) revealing submicron features in the actin (orange arrow). Bottom row: RDM similarly resolves actin filaments and mitochondria where deconvolution does not.
2.4 High NA multicolor fluorescence microscopy
@@
Our second experimental system to demonstrate RDM is a high magnification, high numerical aperture (NA) microscope. Such devices are critical to observing biological samples at subcellular resolution. However, as the combination of magnification and NA increases, so do field-varying aberrations. RDM offers a pathway to utilize the level of magnification and NA needed for subcellular imaging whilst maintaining isotropic resolution over the entire FoV. Moreover, RDM does this efficiently over multiple fluorescence color channels, allowing for multicolor, subcellular-resolution imaging over the entire FoV. To demonstrate this, we image the fluorescence labeled actin (green channel) and mitochondria (red channel) of Bovine Pulmonary Artery Endothelial (BPAE) cells with a 100x magnification, 1.4 NA objective. More details of the sample and imaging tools are found in 4.5.
@@
As per the RDM pipeline, we first calibrate this objective with a single image of randomly scattered beads (see Fig. 3 a)). There are two options for calibrating RDM in the multicolor setting. The first option is to perform a separate calibration for each color channel of interest, with different bead images corresponding to the the different emission wavelengths; this strategy allows RDM to additionally correct chromatic aberrations. The simpler option is to measure only a single set of beads (one image), obtain one set of Seidel coefficients, then generate the synthetic PSFs for each channel from these coefficients according to the desired wavelength. The latter assumes no chromatic aberration and only accounts for wavelength-dependent diffraction effects. Thus, we opt for the former option, though there is very little difference between the resulting estimated red and green Seidel coefficients.

### v5

# bibliography.bib

@@ Abstract
The most ubiquitous form of computational aberration correction for microscopy is deconvolution. However, deconvolution relies on the assumption that the point spread function is the same across the entire field-of-view. This assumption is often inadequate, but space-variant deblurring techniques generally require impractical amounts of calibration and computation. We present a new imaging pipeline that leverages symmetry to provide simple and fast spatially-varying aberration correction. Our ring deconvolution microscopy (RDM) method leverages the rotational symmetry of most microscopes and cameras, and naturally extends to sheet deconvolution in the case of lateral symmetry. We formally derive theory and algorithms for image recovery and additionally propose a neural network based on Seidel coefficients as a fast alternative, as well as extension of RDM to blind deconvolution. We demonstrate significant improvements in speed and image quality as compared to standard deconvolution and existing spatially-varying deconvolution across a diverse range of microscope modalities, including miniature microscopy, multicolor fluorescence microscopy, point-scanning multimode fiber micro-endoscopy, and light-sheet fluorescence microscopy. Our approach enables near-isotropic, subcellular resolution in each of these applications.
@@
Figure 1: Ring Deconvolution Microscopy (RDM). a) Point sources at the sample plane (left) are imaged (right) to point spread functions (PSFs) with a rotationally-symmetric imaging system. The PSFs are linear revolution-invariant (LRI)—they vary with distance from the center of the field-of-view (FoV) (top row), but maintain the same shape at a fixed radius $r$, just revolved around the center (bottom row). b) The RDM pipeline. A one-time calibration procedure (top) captures a single image of randomly-placed point sources (e.g., fluorescent beads), then fits the primary Seidel coefficients (see Sec. 4.2). Next, we either use the Seidel coefficients to generate a radial line of synthetic PSFs, if using ring deconvolution, or we feed the coefficients directly into Deep Ring Deconvolution (DeepRD). After calibration, we can deblur images (bottom) using either ring deconvolution or DeepRD. c) Experimental deblurring of live tardigrade samples imaged with the UCLA Miniscope [aharoni]. From left to right: measurement, standard deconvolution, ring deconvolution, and DeepRD. Ring deconvolution and DeepRD consistently outperform deconvolution.
@@ 1 Introduction
Much of optical engineering is focused on reducing aberrations by adding additional corrective optical elements to an imaging system—consider a microscope objective, comprised of numerous lenses stacked in a housing. Such designs allow for high-performance imaging, but incur added cost, weight, and complication. Even with large and expensive lens stacks, it is difficult and, in some cases, impossible to correct all aberrations across a large area, and so aberrations are often what limit the usable field-of-view (FoV) of a system. Furthermore, some systems cannot accommodate any aberration-correction optics; for example, additional elements may not fit in miniaturized microscopes (e.g. Miniscope [aharoni]) and are prohibitively expensive for large-aperture telescopes [stahl2009preliminary].
@@
Faced with a poorly-corrected imaging system, the modern microscopist instead turns to computational aberration correction, where the burden is shifted onto computer algorithms applied after the image is acquired. The most commonly used correction technique, image deconvolution, captures a calibration image of a small point-like source, known as the point spread function (PSF), in order to characterize the aberrations. The PSF can then be used to computationally deconvolve any image taken with the system via simple and fast algorithms, to yield a deblurred result. A main limitation of this approach is that it assumes that the system’s PSF does not vary spatially (i.e. the system is linear space-invariant, LSI). This assumption is usually only true near the center of the FoV, and optical designers often artificially sacrifice part of the system’s FoV in order to maintain space invariance.
@@
To go beyond space-invariant limitations, a large community effort has gone towards heuristic forms of spatially-varying ’deconvolution’, wherein one measures PSFs at multiple points within the FoV and uses them to correct the image. Such heuristics include: assuming each region of an image is locally LSI [section-Trussell], adaptively splitting the FoV by first quantifying the degree of space-variance [quant-Lin, quant-lohmann], interpolating PSFs [interp-denis, interp-Nagy], decomposing the PSF into space-invariant orthogonal modes [modes-denis, modes-flicker, modes-Hong, modes-miraut, modes-popkin, modes-Sroubek, modes-todlauer, Turcotte:20], and doing the same in Fourier space [modes-fourier-Lee]. These heuristics can approach rigorous recovery as the number of PSFs collected grows, possibly into the hundreds of thousands; however, the tradeoff in terms of the complexity of calibration and computation quickly becomes intractable. For example, in patch-wise deconvolution, the FoV is divided into patches, each of which gets deconvolved by a PSF measured at its center. Maximum accuracy is achieved when the patch size is reduced to a single pixel, but then a megapixel image would require a million PSF measurements and a computation time of hundreds of hours to deblur (see Fig. 5c).
@@
Another emerging modality is deep deblurring [Yanny-multiwiener, ren2018deep, Deng:22, dong2020deep, weigert2018content, UNet_Deblur, li2022incorporating], in which varying amounts of system information are incorporated into a deep neural network. On one end of the spectrum, networks that are primarily data-driven struggle with extrapolation beyond the training data, and tend to reproduce whatever biases existed therein, a particularly relevant point since many of them are trained on simulated data. On the other end, networks that incorporate physical information, such as calibrated PSFs, may have better generalization properties but suffer the same accuracy/efficiency tradeoff as patch-based methods. For these reasons, spatially-varying deblurring has not become commonplace among practitioners. Thus, there is a need for spatially-varying deblurring methods that are effective, efficient, and robust.
Here, we propose a spatially-varying method that requires only a single calibration image and has reasonable compute time (tens of seconds on a megapixel image), while offering rigorous deblurring for imaging systems that are symmetric in some way. We focus on rotationally-symmetric systems—i.e., systems that are symmetric about their optical axis—but also show an example with the lateral symmetry present in light-sheet microscopy. Rotational symmetry occurs in many imaging systems by design, and a significant portion of optical theory is developed under this assumption. While some existing deblurring techniques have leveraged rotational symmetry, they are approximate and restricted to a specific subset of radially-varying blurs: those due to camera zoom [Webster_motion_blur, boracchi2008deblurring], the specific case of a parabolic mirror [luan2018suppress], and an approximate scheme only for blurs from a single lens by applying deconvolution to four concentric regions [zhang_radial_blur, zhang_restoration, zhang_analysis, Zhang_ringing] . Other work does the same for DSLR cameras and also requires three image channels from an RGB camera [inproceedings] . In contrast, what we propose applies to any rotationally-symmetric imaging system, can incorporate more complex PSFs—even if they cannot be theoretically derived, makes no approximations (e.g., isoplanatic regions) in the image formation model, and can easily extend to other symmetries.
@@
Our Ring Deconvolution Microscopy (RDM) technique achieves both accuracy and efficiency. RDM models image formation for rotationally-symmetric imaging systems rigorously, allowing for accurate deblurring while still remaining practical, both computationally and in terms of calibration. The first step in RDM is a simple, single-shot calibration scheme, in which the system’s primary Seidel aberration coefficients are estimated from a single image of randomly distributed point sources. These coefficients quantify the severity of spatial variance and provide the necessary system information for the second step, deblurring. Within this second step we propose two alternative image deblurring algorithms. The first (our main algorithm) is ring deconvolution, which uses a new and rigorous theory for rotationally-symmetric imaging to deblur the image at all points in the FoV, with only order $N^{3}\log(N)$ ($N$ is the image side length) compute time, as compared to $N^{4}$ for full spatially-varying deblurring. For even faster computation (but without theoretical guarantees), we further propose an alternate neural network-based algorithm called Deep Ring Deconvolution (DeepRD), which constrains learning with physical knowledge provided by the system’s Seidel aberration coefficients. If the system is not spatially-varying or minimally so, RDM still offers an improvement over standard deconvolution by instead deconvolving with a synthetic PSF generated by the Seidel coefficients (see Sec. 2.5).
Although ring deconvolution is specific to systems exhibiting rotational symmetry, our theory can be easily adapted to exploit other forms of symmetry. As an example, we derive an analogous form of deconvolution for situations where the blur varies laterally (along one Cartesian axis), which we term sheet deconvolution. This is the case in light-sheet microscopy, where the light-sheet illumination causes space-varying blur in the direction perpendicular to the imaging plane. We show experimental results for sheet deconvolution on light-sheet fluorescence microscopy data in Sec. 4.4, in which we effectively increase the FoV by about twofold through our deblurring process.
@@
These algorithms outperform existing methods, approaching subcellular, isotropic resolution across the FoV. We demonstrate this on four diverse microscope modalities: miniature microscopy, multicolor fluorescence microscopy, multimode fiber micro-endoscopy, and light-sheet fluorescence microscopy. Each of these modalities contains different characteristics and imaging mechanisms that are representative of a wide range of imaging systems, thereby forming a comprehensive basis to demonstrate the wide applicability and practical relevance of our methods. See Fig. 1 for a summary of the RDM pipeline along with an example result on images of live tardigrades. An open source implementation of RDM and its extensions can be found here.
@@ 2 Results
Before we display our experimental results, we briefly outline the RDM pipeline from calibration to deblurring. Further details about RDM can be found in Sec. 4.
2.1 Ring Deconvolution Microscopy Pipeline
The first step in our RDM pipeline is calibration; we measure the system’s response to a point source, i.e., its PSF. The PSFs of a space-varying system will vary across the FoV, and many PSF measurements may be required to fully characterize the system. However, space-varying systems that are rotationally symmetric require fewer measurements for system characterization, which we exploit here. Intuitively, PSFs that are the same distance from the center of the FoV all have the same shape, just rotated at different angles (see Figure 1a) because of the symmetry. We call this property of the PSFs linear revolution-invariance (LRI), and denote it mathematically as
@@ 2.1 Ring Deconvolution Microscopy Pipeline
where $\tilde{h}(\rho,\phi;r,\theta)$ is the (spatially-varying) PSF in polar coordinates from a point source at location $(r,\theta)$. Note that the shape of the PSF itself is not necessarily rotationally symmetric. LRI greatly improves complexity of the calibration procedure, in that we need only measure the PSF at any one point along the circle for each radius $r$ from the optical center.
In practice, directly measuring the PSF at every radius $r$ is impractical. Instead, under LRI, there is a simple and effective method for simultaneously estimating these PSFs from a single image, without any motion stage required. The method is described in Section 4.2 and is the first step of the RDM pipeline. It works by estimating the Seidel aberration coefficients from an image of point sources randomly scattered across the FoV (see Fig. 1b). Seidel coefficients are a natural choice for LRI systems, since Seidel polynomials are explicit functions of the field position and thus use the same coefficients regardless of the position in the FoV [born2013principles]. Thus, they are easier to estimate than Zernike coefficients, which are different at each field position [Zheng:13, Shao:19]. Our experiments demonstrate that the 5 primary (4th order) Seidel coefficients (sphere, coma, astigmatism, field curvature, and distortion) suffice to characterize spatially-varying LRI systems. A discussion of higher-order aberrations can be found in Appendix D.1.
After calibration, the next step is to use the estimated Seidel coefficients to deblur the measured image. LRI systems, much like LSI systems, allow for computationally-efficient models. For LSI systems, we leverage two axes of space-invariance to model the blur as a 2D convolution, and for LRI systems we can leverage the revolution-invariant angular axis, leading to a blur computation that is a sum of 1D convolutions. Section 4.1 provides a description of the forward and inverse algorithms, which we call ring convolution and ring deconvolution, respectively. Therein is also included a theoretical proof of ring convolution’s exactness under rotational symmetry. We briefly describe ring deconvolution here, along with an alternate method using deep learning, termed DeepRD. These methods form the second step in the RDM pipeline (Fig. 1b).
@@
    Ring deconvolution. We derive an optimal algorithm for reconstructing the underlying sample from a blurry image given the PSF at each radius of the FoV of a rotationally-symmetric imaging system. This is our main algorithm.
@@
    Deep Ring Deconvolution (DeepRD). Although ring deconvolution is significantly faster to compute than a full patch-based spatially-varying deblur technique, it may still be relatively slow (on the order of a few minutes) for very large image sizes (e.g., beyond 1024×1024)1024$\times$ 1024)1024 × 1024 ) or video data. Deep learning enables a faster (but approximate) version of ring deconvolution called DeepRD. As input, it takes a blurry image and a list of the 5 primary Seidel coefficients. DeepRD is trained on a dataset of natural images that are synthetically blurred using ring convolution.
@@
Figure 2: Ring Deconvolution Microscopy (RDM) for the UCLA Miniscope. After calibrating the Miniscope with a single image of fluorescent beads (see Fig. 1b), we show results from several deblurring methods for comparison: standard deconvolution (using a single measured PSF), a U-Net trained on our spatially-varying blur dataset, ring deconvolution and DeepRD. Deconvolution assumes space-invariance while the remaining methods are designed to handle spatially-varying aberrations. In the first row, ring deconvolution and DeepRD clearly resolve resolution target elements near the edges of the FoV, which are not as well resolved by the two other methods. Zoom-ins show RDM resolves up to element 6 of group 9 (blue inset) and element 5 of group 8 (green inset). Similar results along with corresponding insets are shown for the other samples.
2.2 Experimental Results
Miniature microscopy
@@ 2.2 Experimental Results
Our first experimental system is the UCLA Miniscope [aharoni], a miniature microscope used largely for neuroimaging in freely-behaving animals. The first step of our experimental pipeline is to calibrate the imaging system by capturing a single image of randomly-placed fluorescent beads (obtained by smearing bead solution on a slide). We use it to fit Seidel coefficients, obtaining 0.85, 0.56, 0.25, 0.29, and 0 waves of spherical aberration, coma, astigmatism, field curvature and distortion, respectively. These numbers, while specific to our particular assembly of the Miniscope, are consistent with the aberration profile of a radial GRIN lens [Wang:90], which is the objective lens used by the Miniscope. The fact that the off-axis coefficients (i.e. all of the primary coefficients except for spherical) are nonzero confirms that the system is indeed spatially-varying. For comparison with standard deconvolution deblurring, we also acquire the center PSF by imaging a single fluorescent bead isolated and centered in the FoV. The PSF is then denoised before its use in deconvolution (see 4.6).
@@
Having completed the one-time calibration for the Miniscope, we image a variety of different samples: a USAF resolution target, rabbit liver tissue, and live fluorescence-stained tardigrades. For each sample, we compare reconstructions from ring deconvolution and DeepRD, as well as standard deconvolution (using the PSF measured at the center of the FoV) and a baseline U-Net (see Fig. 2).
@@
The USAF resolution target was placed at 9 separate locations in the FoV, which allows us to observe the smallest features at each location. We stitch these 9 images together, using only the region of each constituent image that contains the high-resolution group, to form the measurement image shown in Fig. 2 (first row). Ring deconvolution and DeepRD—having knowledge of the field-varying aberrations via the Seidel coefficients—give the most improvement near the edges and corners of the image. Standard deconvolution produces a noisy, low contrast result in those regions due to the mismatch of the center PSF with the edge PSFs. We also note that both learning methods (U-Net and DeepRD) offer the best denoising performance, a well-known property of neural networks [vincent2008extracting]; however, this comes at the cost of inconsistent performance—both models perform worse on the resolution target than on the other samples.
@@
For our second sample, rabbit liver tissue, our methods perform well, revealing features in the corners of the image, including the outlines of membranes, which are not clear in the raw measurement or other methods. Such clarity is critical for downstream tasks like image segmentation.
@@
Finally, we fluorescently stain live semi-starved tardigrades with a DNA gel stain and capture a series of videos. We apply deblurring to each frame and display one such frame in the bottom row of Fig. 2; the full videos can be found here. As with the previous experiments, ring deconvolution and DeepRD provide increased image contrast and detail in the corners of the frames compared to the other methods. In particular, the small, dot-like features within the tardigrade are better resolved.
@@
Further experimental details can be found in Sec. 4.6. The baseline U-Net we use is inspired by the popular Content-Aware Image Restoration (CARE) model [weigert2018content]; we use a similar model structure but train both DeepRD and the baseline U-Net using a novel synthetic training strategy: applying ring convolutions using PSFs generated from random Seidel coefficients to natural images from Div2k [agustsson2017ntire](see Sec. 4.3). The dataset is publicly available here. Despite being trained on simulated data, both the U-Net and DeepRD exhibit reasonable deconvolution performance on experimental images, as shown in Fig. 5. Further, in Appendix E, we provide a demonstration that DeepRD is interpretable: it understands the individual effects of each aberration coefficient.
High NA multicolor fluorescence microscopy
@@
Our second experimental system to demonstrate RDM is a high magnification, high numerical aperture (NA) microscope. Such devices are critical to observing biological samples at subcellular resolution. However, as the NA increases, so do field-varying aberrations. RDM offers a pathway to utilize the level of magnification and NA needed for subcellular imaging whilst maintaining isotropic resolution over the entire FoV. Moreover, RDM does this efficiently over multiple fluorescence color channels, allowing for multicolor, subcellular-resolution imaging over the entire FoV. To demonstrate this, we image the fluorescently-labeled actin (green channel) and mitochondria (red channel) of Bovine Pulmonary Artery Endothelial (BPAE) cells with a 100$\times$ 1.4 NA objective. More details of the sample and imaging tools are found in 4.6.
@@
Again, we calibrate the system by capturing a single image of randomly scattered beads (see Fig. 3 a)). We choose to perform a separate calibration for each color channel, with different bead images corresponding to the different emission wavelengths; this strategy allows RDM to additionally correct chromatic aberrations. In practice, however, there is very little difference between the resulting estimated red and green Seidel coefficients.